# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ Abstract
Data freshness, measured by Age of Information (AoI), is highly relevant in networked applications such as Vehicle to Everything (V2X), smart health systems, and Industrial Internet of Things (IIoT). Yet, freshness alone does not equate to informativeness. In decision-critical settings, some stale data may prove more valuable than fresh updates. To explore this nuance, we move beyond AoI-centric policies and investigate how data staleness impacts decision-making under data-staleness-induced uncertainty. We pose a central question: What is the value of information, when freshness fades—and only its power to shape remote decisions remains? To capture this endured value, we propose AR-MDP, an Age-aware Remote Markov Decision Process framework, which co-designs optimal sampling and remote decision-making under a sampling frequency constraint and random delay. To efficiently solve this problem, we design a new two-stage hierarchical algorithm namely Quick Bellman-Linear-Program (QuickBLP), where the first stage involves solving the Dinkelbach root of a Bellman variant and the second stage involves solving a streamlined linear program (LP). For the tricky first stage, we propose a new One-layer Primal-Dinkelbach Synchronous Iteration (OnePDSI) method, which overcomes the re-convergence and non-expansive divergence present in existing per-sample multi-layer algorithms. Through rigorous convergence analysis of our proposed algorithms, we establish that the worst-case optimality gap in OnePDSI exhibits exponential decay with respect to iteration $K$ at a rate of $\mathcal{O}(\frac{1}{R^{K}})$. Through sensitivity analysis, we derive a threshold for the sampling frequency, beyond which additional sampling does not yield further gains in decision-making. Simulation results validate our analyses.
@@ I Introduction
A significant challenge in the field is the scarcity of rigorous analytical frameworks that establish clear relationships between information freshness and value. In many scenarios, freshness alone does not determine how valuable a data update is to the eventual decision it informs. Instead, the value of information may depend on multiple interrelated factors beyond AoI, including the semantic content of transmitted packets and the underlying dynamics of the monitored source[40, 41, 42, 43]. This recognition has driven the development of various AoI variants. One approach introduces nonlinear AoI penalties, implemented through both empirical configurations [44, 45, 46, 47] and theoretically derived functions [48, 49, 50, 51]. These nonlinear formulations aim to quantify the loss resulting from information staleness. Additionally, researchers have proposed various heuristic metrics, including Age of Synchronization (AoS) [52], Age of Incorrect Information (AoII) [53], and Age of Changed Information (AoCI) [54], and Age of Collected Information [55, 56]. These metrics customize time-related penalization from a wider perspective than what can be captured with age, particularly in applications involving rapidly evolving source dynamics. For remote estimation, mean square estimation error (MSEE) [16] and context-aware Urgency of Information (UoI) [57] are leveraged to penalize the real-time reconstruction distortion. Although information freshness is widely regarded as beneficial in communication systems, its specific value to decision quality has not been fully characterized. This leads to a key question: From the perspective of remote decision-making, what kind of information truly matters, and how should its value guide what is sampled and transmitted? This open question motivates our theoretical investigation in this paper.
@@
Several works have proposed heuristic approaches to characterize this relationship. In [58], Dong et al. introduced Age upon Decisions (AuD), which measures the time elapsed between data generation and its use in decision-making, where the decision epoch follows a stochastic distribution. In a similar vein, [59] proposed Age of Actuation (AoA). In [60], three types of decisions: correct decisions, incorrect decisions, and missed decisions are assigned different time-cost functions. A new metric termed Penalty upon Decision (PuD) was proposed. In [41] and [61], a tensor-based metric termed Goal-oriented Tensor (GoT) was proposed as a unified framework for existing metrics. However, while prior work has advanced communication optimization through a variety of metrics, it often overlooks how decision-making systems actually function when operating with potentially stale information. The core challenge is to bridge the gap between metric-driven optimization and the realities of decision-making under data-staleness-induced uncertainty.
@@
In this paper, we aim to examine how stale information impacts remote stochastic decision-making under conditions of data-staleness-induced uncertainty. To this end, we propose the Age-aware Remote Markov Decision Process [^1] In [62], the term remote MDP was first proposed as a pathway to pragmatic or goal-oriented communications. Our paper focuses on the communication delay and introduces the age to enhance remote decision-making to achieve a certain goal, hence the term age-aware remote MDP. (AR-MDP), a comprehensive framework that jointly optimizes sampling and sequential decision-making, with a specific purpose to achieve goal-oriented effective decisions. The relationship between sampling and decision-making exhibits inherent bidirectional coupling. Sampling policies directly influence the staleness-induced uncertainty, which can lead to unsatisfactory decision outcomes. Conversely, decision-making processes affect the stochastic evolution of the source system, which in turn impacts the effectiveness of content-driven goal-oriented sampling mechanisms. To decouple these two processes and achieve optimal decision making under random delay and sampling frequency constraint, we formulate the problem as a constrained Partially Observable Semi-Markov Decision Process problem, where AoI no longer serves as a typical indicator, but instead plays a critical role as side information for enhancing remote, goal-oriented decisions.. We design efficient algorithms to solve this problem. To the best of our knowledge, this is the first work that introduces AoI into the time-lag MDP family, and the first work that explores AoI’s new role as side information to facilitate remote decision making under random delay.
@@ II-A Sampling Under Random Delay
The results in this paper contribute to the optimal sampling design under random delay. In real-world network environments, communication channels inevitably experience random delays due to various factors: network handover, congestion, variable sample sizes, and packet retransmissions. These fundamental characteristics have driven research into developing optimal sampling policies under random delay. Existing literature has focused on optimizing three key aspects: $i$) information freshness [63, 64, 65, 66, 67, 68]; $ii$) remote estimation [25, 16, 17, 69, 70, 71]; and $iii$) remote inference [72, 73, 74, 75] under random delay. A particularly noteworthy and counter-intuitive finding in this field reveals that optimal sampling may require the source to deliberately wait before submitting a new sample to the channel, challenging the conventional wisdom of throughput-optimal zero-wait sampling policy.
@@
Information Freshness: In the seminal work [63], Sun et al. derived an AoI-optimal sampling policy under random delay. This paper revealed that under a maximum rate constraint, the AoI-optimal sampling follows a threshold structure, where sampling is activated only when the current AoI exceeds a specific threshold determinable through a low-complexity bisection search method. In [64], the optimal sampling policy for a non-linear monotonic function of AoI was designed. Tang et al. [65] extended this framework to scenarios with unknown delay statistics, employing stochastic approximation methods to determine the AoI-optimal sampling policy under unknown delay statistics. Further advancements were made by Pan et al. [66], who developed AoI-optimal sampling policies under unreliable transmission with random two-way delay. Most recently, Liyanaarachchi and Ulukus enhanced [63] by incorporating random ACK delay, demonstrating that sampling before receiving acknowledgment can potentially achieve superior AoI performance [67]. In [68], Peng et al. designed optimal sampling policies that achieve minimal Age of Changed Information (AoCI) [76]—a metric capable of detecting source changes—under known and unknown delay statistics respectively.
@@
Remote Estimation: The theoretical foundations of remote estimation-oriented sampling under random delay were established through [16, 17]. These studies developed an optimal sampling policy for the Wiener process that minimizes the mean square estimation error (MSEE) while adhering to sampling frequency constraints. Their research revealed that the optimal sampling policy exhibits a threshold structure, where sampling is initiated only when the real-time MSEE surpasses a predetermined threshold. Building upon this foundation, Ornee et al. [25] expanded this theoretical framework by investigating MSEE-optimal sampling for the Ornstein-Uhlenbeck (OU) process, a stationary Gauss-Markov process, under random delay. In [69] and [70], the MSEE-optimal sampling policies are derived for the Wiener process and the OU process under unknown delay statistics. In [71], Chen et al. derived the optimal sampling policy that achieves minimized uncertainty-of-information under random delay, where UoI is defined as the conditional entropy of the source at the receiver given the observation history [51]—mathematically expressed as $H(X_{t}|\mathcal{H}_{t})$, with $X(t)$ representing the source state at time $t$ and $\mathcal{H}_{t}$ denoting the available observation history at the receiver. Most recently, Chen et al. [77] derived AoII-optimal sampling policies under random delay.
@@
Remote Inference: Recent research has revealed insights into remote inference performance and its relationship with information freshness metrics. In [73, 72], Shisher et al. demonstrated that the loss function in remote inference may not be monotonic in terms of the age of the samples (features) used, if the source sequence is not Markovian. Upon making this remarkable observation, the authors developed policies that allow selection of aged samples from the buffer, rather than the freshest one. This was termed the “selection-from-buffer” model. In [74], a learning and communication co-design problem that jointly optimizes feature length selection and transmission scheduling is proposed. In [75], Ari et al expanded previous works by incorporating time-varying statistics of random delay and delayed feedback, developing optimal sampling policies to minimize long-term inference error within the “selection-from-buffer” model.
@@ II-B Decision-Making over Stale Status
Figure 2: Evolution of the age $\Delta(t)$ over time.
However, the above time-lag MDPs, where the observation delay follows a given distribution (DDMDP can be regarded as a special type of SDMDP), potentially assume that the state is sampled and transmitted to the decision maker at every time slot [^2] In this case, each state $X_{i},\forall i\in\{0,1,...n\}$ are all sampled and forwarded to the decision maker. The observation delay $D$ is i.i.d random variable and is independent of the sampling policy.. This setup presumes that the system can transmit every state update without encountering any backlog. In practice, constantly sampling and transmitting may result in infinitely accumulated packets in the queue, resulting in severe congestion. This motivates the need for queue control and adaptive sampling policy design in the network [81, 63, 22, 65, 82, 83], where Age of Information (AoI) serves as a key performance indicator. Suppose the $i$-th sample is generated at time $S_{i}$ and is delivered at the receiver at time $D_{i}$. Then, AoI is defined as a sawtooth piecewise function:
$\Delta(t)=t-S_{i},D_{i}\leq t<D_{i+1},\forall i\in\mathbb{N},$ (1)
as shown in Fig. 2. From this definition, the most recently available information at the receiver at time slot $t$ is $O(t)=X_{t-\Delta(t)}$. Different from the DDMDP and SDMDP where the time lag is a constant $d$ or an i.i.d random variable $D$, with $O(t)=X_{t-d}$ or $O(t)=X_{t-D}$, the time lag in the practical network with queue and finite transmission rate is a controlled stochastic process $\Delta(t)$, with $O(t)=X_{t-\Delta(t)}$.
To the best of our knowledge, the design of optimal remote decision making under controlled stochastic delay remains an unexplored research direction, which we address in this paper by studying AR-MDP.
@@ II-C The Novelty of Our Work
  • System Model: This paper proposes AR-MDP, a novel theoretical framework integrating optimal sampling and decision-making under random delay. Differing from prior sampling designs which often treat information as an end in itself—optimizing for freshness, accuracy, or estimation quality under random delays [63, 64, 65, 66, 67, 68, 72, 73, 74, 75, 25, 16, 17, 69, 70, 71], we treat information as a means to action, where its value is defined not by how precise it is, but by how well it enables timely and effective decisions under uncertainty and delay. By modeling the controlled stochastic “delay” as the AoI $\Delta(t)$, AR-MDP also generalizes classical time-lag MDP theoretical framework [79, 80] into a practically relevant model. Through sensitivity analysis, we uncover a fundamental sampling frequency threshold. Beyond this threshold, further increasing the sampling frequency yields no improvement in decision performance. This finding reveals an information-theoretic limit: in decision-centric systems under random delay, the value of information is not solely determined by its freshness, but by its timely relevance to control.
@@
  • Methodology Design: We design QuickBLP, a computationally efficient single-layer algorithm that addresses limitations associated with iterative re-convergence common in multi-layer per-sample algorithms, e.g., [71, Algorithm 1], [84, Algorithm 1], and [85, Section IV.C]. This algorithm is designed based on a key analytical insight that the optimal solution exhibits a threshold structure and can be obtained through a two-stage process. The first stage determines the Dinkelbach root of a Bellman variant, for which we develop OnePDSI, a Cauchy sequence that converges asymptotically to the root without requiring re-convergence. The second stage involves finding the Dinkelbach root of a per-sample constrained Markov Decision Process (CMDP), which traditionally necessitates multiple CMDP solutions and suffers from re-convergence. We resolve this challenge by proving that the Dinkelbach root can be explicitly calculated through the optimal value of a linear program (LP), enabling direct solution to the root through a single LP solution. To the best of our knowledge, QuickBLP is the first framework to tackle constrained partially observable SMDPs through a streamlined single-layer flow, fundamentally improving efficiency by eliminating re-convergence overhead.
@@
  • Theoretical Rigor and Convergence: We significantly advance our previous work [1] by resolving critical convergence limitations. To overcome inherent divergence risks, we design two novel algorithms in this paper: Bisec-$\tau$-RVI and OnePDSI. Although convergence analyses often rely on the Banach Contraction Mapping Theorem [86, Theorem 6] [66, 85], this method falls short in capturing the behavior of our models. Our approach departs from this tradition, providing rigorous proofs that both algorithms guarantee efficient exponential convergence in worst-case settings. These guarantees reinforce the reliability and efficiency of our algorithms in practical applications.

### v3

@@ Abstract
Data freshness, measured by Age of Information (AoI), is highly relevant in networked applications such as Vehicle to Everything (V2X), smart health systems, and Industrial Internet of Things (IIoT). However, freshness alone does not always equate to utility in decision-making. In decision-critical settings, some stale data may be more valuable than fresh updates. Motivated by this, we move beyond AoI-centric policies and investigate how data staleness affects remote decision-making effectiveness under random delay and limited communication resources. To this end, we propose AR-MDP, an Age-aware Remote Markov Decision Process framework, which co-designs optimal sampling and remote decision-making under a sampling frequency constraint and random delay. To efficiently solve this problem, we design a new two-stage hierarchical algorithm, namely Quick Bellman-Linear-Program (QuickBLP), where the first stage involves solving the Dinkelbach root of a Bellman variant and the second stage involves solving a streamlined linear program (LP). For the tricky first stage, we propose a new One-layer Primal-Dinkelbach Synchronous Iteration (OnePDSI) method, which overcomes the re-convergence and non-expansive divergence present in existing per-sample multi-layer algorithms. Through rigorous convergence analysis of our proposed algorithms, we establish that the worst-case optimality gap in OnePDSI exhibits exponential decay with respect to iteration $K$ at a rate of $\mathcal{O}(\frac{1}{R^{K}})$. Through sensitivity analysis, we derive a threshold for the sampling frequency, beyond which additional sampling does not yield further gains in decision-making. Simulation results validate our analyses.
@@ I Introduction
A significant challenge in the field lies in the lack of unified analytical frameworks that link information freshness to its effectiveness in real-time decision-making. In many scenarios, freshness alone does not determine how beneficial an update is to the downstream decision-making task. Instead, the effectiveness of information may depend on multiple interrelated factors beyond AoI, including the semantic content of transmitted packets and the underlying dynamics of the monitored source[97, 50, 35, 77]. This recognition has driven the development of various AoI variants. One approach introduces nonlinear AoI penalties, implemented through both empirical configurations [84, 45, 23, 6] and theoretically derived functions [95, 82, 101, 18]. These nonlinear formulations aim to quantify the loss resulting from information staleness. Additionally, researchers have proposed and optimized various heuristic metrics, including Age of Synchronization (AoS) [108], Age of Incorrect Information (AoII) [59, 25, 22], Age of Changed Information (AoCI) [100], and Age of Collected Information [98, 37]. These metrics customize time-related penalization from a wider perspective than what can be captured with age, particularly in applications involving rapidly evolving source dynamics. For remote estimation, mean square estimation error (MSEE) [87] and context-aware Urgency of Information (UoI) [107] are leveraged to penalize the real-time reconstruction distortion. Despite these advances, the relationship between AoI and decision-making performance is not fully characterized. Specifically, existing studies often optimize communication metrics (e.g., AoI or AoII) without explicitly modeling how delayed information influences sequential decision-making outcomes. This motivates the question: How does delayed or outdated information affect the quality of remote decisions, and how should communication policies adapt accordingly?
@@
Several works have proposed heuristic approaches to characterize this relationship. In [30], Dong et al. introduced Age upon Decisions (AuD), which measures the time elapsed between data generation and its use in decision-making, where the decision epoch follows a stochastic distribution. In a similar vein, [66] proposed Age of Actuation (AoA). In [46, 74, 78], Cost of Actuation Error (CoAE) was proposed to penalize distortion-induced error actuation. In this setting, a penalty $C_{i,j}$ is incurred when the true system state is $i$ while the remote controller makes decisions based on an estimated value $\hat{X}_{t}=j$. This line of work primarily focuses on the estimation of a discrete-time Markov chain (DTMC), and quantifies the semantic mismatch between state and inferred action due to delayed or lossy communication. In [111], three types of decisions: correct decisions, incorrect decisions, and missed decisions are assigned different time-cost functions. A new metric termed Penalty upon Decision (PuD) was proposed. In [50] and [51], a tensor-based metric termed Goal-oriented Tensor (GoT) was proposed as a unified framework for existing metrics. However, while prior work has advanced communication optimization through a variety of metrics, it often overlooks how decision-making systems actually function when operating with potentially stale information. Bridging this gap requires a framework that explicitly models how stochastic queuing delays influence decision quality in dynamic systems.
@@
In this paper, we aim to examine how stale information impacts remote stochastic decision-making under communication constraints. To this end, we propose the Age-aware Remote Markov Decision Process [^1] In [34], the term remote MDP was first proposed as a pathway to pragmatic or goal-oriented communications. Our paper focuses on the communication delay and introduces the age to enhance remote decision-making to achieve a certain goal, hence the term age-aware remote MDP. (AR-MDP), a comprehensive framework that jointly optimizes sampling and sequential decision-making, with a specific purpose to achieve goal-oriented effective decisions. The relationship between sampling and decision-making exhibits inherent bidirectional coupling. Sampling decisions affect the freshness of information available for the remote decision maker, which can result in unsatisfactory decision outcomes. Conversely, decision-making processes affect the stochastic evolution of the source system, which in turn impacts the effectiveness of content-driven goal-oriented sampling mechanisms. To decouple these two processes and achieve optimal decision-making under random delay and a sampling frequency constraint, we formulate the problem as a constrained partially observable semi-Markov decision process, where AoI no longer serves as a typical indicator, but as side information that informs delay-aware decision-making. We design efficient algorithms to solve this problem. To the best of our knowledge, this is among the first attempts to treat AoI as dynamic side information in remote decision-making systems, and to systematically integrate it into a formal decision-theoretic framework.
@@ II-A Sampling Under Random Delay
The results in this paper contribute to the optimal sampling design under random delay. In real-world network environments, communication channels inevitably experience random delays due to various factors: network handover, congestion, variable sample sizes, and packet retransmissions. These fundamental characteristics have driven research into developing optimal sampling policies under random delay. Existing literature has focused on optimizing three key aspects: $i$) information freshness [88, 83, 91, 71, 54, 75]; $ii$) remote estimation [68, 87, 86, 93, 21, 19]; and iiiiii) remote inference [80, 81, 79, 4] under random delay. A particularly noteworthy and counter-intuitive finding in this field reveals that optimal sampling may require the source to deliberately wait before submitting a new sample to the channel, challenging the conventional wisdom of throughput-optimal zero-wait sampling policy.
@@
Information Freshness: In the seminal work [88], Sun et al. derived an AoI-optimal sampling policy under random delay. This paper revealed that under a maximum rate constraint, the AoI-optimal sampling follows a threshold structure, where sampling is activated only when the current AoI exceeds a specific threshold determinable through a low-complexity bisection search method. In [83], the optimal sampling policy for a non-linear monotonic function of AoI was designed. Tang et al. [91] extended this framework to scenarios with unknown delay statistics, employing stochastic approximation methods to determine the AoI-optimal sampling policy under unknown delay statistics. Further advancements were made by Pan et al. [71], who developed AoI-optimal sampling policies under unreliable transmission with random two-way delay. Most recently, Liyanaarachchi and Ulukus extended [88] by incorporating random ACK delay, demonstrating that sampling before receiving acknowledgment can potentially achieve superior AoI performance [54]. In [75], Peng et al. designed optimal sampling policies that achieve minimal Age of Changed Information (AoCI) [99]—a metric capable of detecting source changes—under known and unknown delay statistics. Most recently, Chen et al. [22] derived AoII-optimal sampling policies under random delay.
@@
Remote Estimation: The theoretical foundations of remote estimation-oriented sampling under random delay were established through [87, 86]. These studies developed an optimal sampling policy for the Wiener process that minimizes the mean square estimation error (MSEE) while adhering to sampling frequency constraints. Their research revealed that the optimal sampling policy exhibits a threshold structure, where sampling is initiated only when the real-time MSEE surpasses a predetermined threshold. Building upon this foundation, Ornee et al. [68] expanded this theoretical framework by investigating MSEE-optimal sampling for the Ornstein-Uhlenbeck (OU) process, a stationary Gauss-Markov process, under random delay. In [93] and [21], the MSEE-optimal sampling policies are derived for the Wiener process and the OU process under unknown delay statistics. In [19], Chen et al. derived the optimal sampling policy that achieves minimum uncertainty of information under random delay, where UoI is defined as the conditional entropy of the source at the receiver given the observation history [18]—mathematically expressed as $H(X_{t}|\mathcal{I}_{t})$, with $X_{t}$ representing the source state at time $t$ and $\mathcal{I}_{t}$ denoting the available observation history at the receiver.
@@
Remote Inference: Recent research has revealed insights into remote inference performance and its relationship with information freshness metrics. In [81, 80], Shisher et al. demonstrated that the loss function in remote inference may not be monotonic in terms of the age of the samples (features) used, if the source sequence is not Markovian. Upon making this remarkable observation, the authors developed policies that allow selection of aged samples from the buffer, rather than the freshest one. This was termed the “selection-from-buffer” model. In [79], a learning and communication co-design problem that jointly optimizes feature length selection and transmission scheduling is proposed. In [4], Ari et al. expanded previous works by incorporating time-varying statistics of random delay and delayed feedback, developing optimal sampling policies to minimize long-term inference error within the “selection-from-buffer” model. All these works reveal that the remote inference utility may not be a monotonic function in terms of AoI. Together, these works demonstrate that in remote inference settings, the utility of a sample is not necessarily a monotonic function of its age.
@@ II-B Decision-Making over Stale Status
Figure 2: AoI evolution in slotted time. The $i$-th sample is generated at $S_{i}$ and delivered at $D_{i}$ with random delay $Y_{i}$. The control is updated only upon delivery and is held constant until the next delivery. Shaded areas indicate action-holding intervals, during which the delivered observation remains $X_{S_{i}}$ for $t\in[D_{i},D_{i+1})$ while the staleness $\Delta(t)$ increases linearly.
However, the above time-lag MDPs, where the observation delay follows a given distribution (DDMDP can be regarded as a special type of SDMDP), potentially assume that the state is sampled and transmitted to the decision maker at every time slot [^2] In this case, each state $X_{i},\forall i\in\{0,1,...n\}$ are all sampled and forwarded to the decision maker. The observation delay $D$ is an i.i.d. random variable and is independent of the sampling policy.. This setup presumes that the system can transmit every state update without encountering any backlog. In practice, constantly sampling and transmitting may result in infinitely accumulated packets in the queue, resulting in severe congestion. This motivates the need for queue control and adaptive sampling policy design in the network [106, 88, 2, 91, 72, 109], where Age of Information (AoI) serves as a key performance indicator. Suppose the $i$-th sample is generated at time $S_{i}$ and is delivered at the receiver at time $D_{i}$. Then, in a time slotted system, AoI is defined as:
$\Delta(t)=t-S_{i},D_{i}\leq t<D_{i+1},\qquad\forall i\in\mathbb{N},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t\in\mathbb{N}},$ (1)
as shown in Fig. 2. From this definition, the most recently available information at the receiver at time slot $t$ is $O(t)=X_{t-\Delta(t)}$. In slotted time, AoI evolves deterministically between deliveries and resets to the realized delivery delay; specifically
[…]
Different from the DDMDP and SDMDP where the time lag is a constant $d$ or an i.i.d. random variable $D$, with $O(t)=X_{t-d}$ or $O(t)=X_{t-D}$, the effective delay in AR-MDP is sampling-dependent (through $S_{i}$) and coupled with random transmission delay, so the time lag is policy-dependent rather than an exogenous fixed/i.i.d. variable [^3] While the AoI evolves deterministically between successful updates (i.e., linearly increases), its reset events depend on both the sampling actions and the stochastic delay process. Thus, it can be viewed as a process indirectly governed by the sampling actions..
To the best of our knowledge, the design of optimal remote decision-making in the presence of sampling-dependent stochastic observation delays, as captured by the AoI process, remains an unexplored research direction, which we address in this paper by studying AR-MDP.
@@ II-C The Novelty of Our Work
  • System Model: This paper proposes AR-MDP, a novel theoretical framework integrating optimal sampling and decision-making under random delay. Differing from prior sampling designs which often treat information as an end in itself—optimizing for freshness, accuracy, or estimation quality under random delays [88, 83, 91, 71, 54, 75, 80, 81, 79, 4, 68, 87, 86, 93, 21, 19], we treat information as a means to action, where its contribution is defined not by how precise it is, but by how well it enables timely and effective decisions. This goes beyond distortion-based formulations such as the Cost of Actuation Error (CoAE) [46, 74, 32, 78], by embedding the staleness-induced impact directly into the sequential decision-making process. Different from classical time-lag MDPs (e.g., DDMDP/SDMDP) that model the delay as an exogenous quantity where either a fixed constant $d$ or a random variable $D$ independent of the sampling policy, our AR-MDP adopts a different remote-decision information structure in which AoI $\Delta(t)$ is an endogenous staleness process shaped jointly by controllable sampling or waiting decisions and random delay $Y_{i}$. Under the sample-and-hold information structure, we establish an exact fixed-dimensional sufficient statistic and an embedded lifted MDP (Lemma 1), enabling tractable average-cost analysis.
@@
  • Methodology Design: We design QuickBLP, a computationally efficient single-layer algorithm that addresses limitations associated with iterative re-convergence common in multi-layer per-sample algorithms, e.g., [19, Algorithm 1], [7, Algorithm 1], and [110, Section IV.C]. This algorithm is designed based on a key analytical insight that the optimal solution exhibits a threshold structure and can be obtained through a two-stage process. The first stage determines the Dinkelbach root of a Bellman variant, for which we develop OnePDSI, a Cauchy sequence that converges asymptotically to the root without requiring re-convergence. The second stage involves finding the Dinkelbach root of a per-sample constrained Markov Decision Process (CMDP), which traditionally necessitates multiple CMDP solutions and suffers from re-convergence. We resolve this challenge by proving that the Dinkelbach root can be explicitly calculated through the optimal value of a linear program (LP), enabling direct solution to the root through a single LP solution. To the best of our knowledge, QuickBLP is the first framework to tackle constrained partially observable SMDPs through a streamlined single-layer flow, fundamentally improving efficiency by eliminating re-convergence overhead.
@@
  • Theoretical Rigor and Convergence: We significantly advance our previous work [49] by resolving critical convergence limitations. To overcome inherent divergence risks, we design two novel algorithms in this paper: Bisec-$\tau$-RVI and OnePDSI. Although convergence analyses often rely on the Banach Contraction Mapping Theorem [5, Theorem 6] [71, 110], this method falls short in capturing the behavior of our models. Our approach departs from this tradition, providing rigorous proofs that both algorithms guarantee efficient exponential convergence in worst-case settings. These guarantees reinforce the reliability and efficiency of our algorithms in practical applications.
II-D Notations
The main notations throughout this paper are summarized in Table II.
TABLE II: Summary of Notations
Symbol Description
$X_{t}$ System state at time slot $t$
$a_{t}$ Response action taken by the decision maker at time $t$
$u_{t}$ Sampling action at time slot $t$
$\pi_{t}$ History-dependent policy
$\phi_{t}$ State-dependent policy
$\boldsymbol{\psi}$ Policy composed by $\phi_{0:\infty}$
$\boldsymbol{\psi}^{\lambda}$ Policy in Problem 4 induced by parameter $\lambda$
$\mathcal{H}_{t}(\cdot)$ Sufficient statistics function
$\Delta(t)$ AoI at time $t$
$S_{i}$ Time slot when the $i$-th sample is taken
$D_{i}$ Time slot when the $i$-th sample is delivered
$Y_{i}$ Random delivery delay between $S_{i}$ and $D_{i}$
$Z_{i}$ Waiting time to sample the $i$-th sample
$A_{i}$ Holding-action taken at the $i$-th epoch
$G_{i}$ Lifted State of Lifted MDP at epoch $i$
$C(x,a)$ Cost function given state $x$ and action $a$
$\lambda$ Dinkelbach parameter
$\mathscr{P}_{\mathrm{MDP}}(\lambda)$ Transformed MDP tuple give parameter $\lambda$
$\phi_{\lambda}^{*}$ Optimal policy of MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$
$U(\lambda)$ Optimal value of Problem 4
$V^{\star}(\cdot;\lambda)$ Optimal relative value function given $\lambda$
${U}_{K}(\lambda),{V}_{K}(\gamma;\lambda)$ RVI values in (25)
$\tilde{U}_{K},\tilde{V}_{K}(\gamma;\lambda)$ $\tau$-RVI values in (28)
$e_{\mathrm{U}}^{(K)}(\lambda)$$e_{\mathrm{V}}^{(K)}(\cdot;\tau,\lambda)$ Relative errors for the $K$-iteration in $\tau$-RVI given $\lambda$
$\kappa$ A parameter for OnePDSI
$e_{\rho}^{(K)}$$e_{\mathrm{W}}^{(K)}(\cdot;\kappa)$ Relative errors for the $K$-iteration in OnePDSI given $\kappa$
$W^{\star}(\cdot)$ Variables in fixed-point equations (38)
$\rho_{K}$$\widetilde{W}_{K}(\cdot)$ OnePDSI values in (43)
$f_{\max}$ Maximum average sampling rate
$H(\lambda;f_{\max})$ Optimal value of Problem 5
$\theta$ Lagrangian multiplier
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})$ Lagrangian function
$\Upsilon(\theta,\lambda;f_{\max})$ Lagrangian dual function
$d(\lambda;f_{\max})$ Optimal value of Problem 6
$Q^{\star}(f_{\max})$ Optimal value of Problem 7
$\theta_{\lambda}^{\star}$ Optimal variable $\theta$ for fixed $\lambda$ in Problem 6
$\mathcal{Q}^{\lambda}$ Long-term average cost given policy $\phi_{\lambda}^{*}$
$\mathcal{F}^{\lambda}$ Long-term average sampling rate given policy $\phi_{\lambda}^{*}$
$\mathcal{F}^{\lambda^{+}}$ Right limit of $\mathcal{F}^{\lambda}$
$\mathcal{F}^{\lambda^{-}}$ Left limit of $\mathcal{F}^{\lambda}$
$f_{\max}^{T}$ Sapling frequency threshold
$\mathbf{P}_{a}$ Transition probability matrix given action $a$
$\mathbf{P}^{n}$ $n$-step transition probability matrix given action $a$
$\mathbf{P}_{i\times j}$ the $(i,j)$-th entry of the transition matrix $\mathbf{P}$
$\rho^{\star}$ Optimal value of Problem 3; Root of $U(\lambda)$
$h^{\star}$ Optimal value of Problem 1; Root of $H(\lambda;f_{\max})$
$\pi_{a}(\cdot)$ Stationary distribution over states under $\mathbf{P}_{a}$